ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

6. Topology

1991 MS Classification: Prieto, C., and M. Aguilar, Instituto de Matematicas, UNAM, Mexico Quasifibrations and Bott periodicity Following a beautiful idea of Dusa McDuff, complex Bott periodicity theorem is proved using the concept of quasifibration and simple linear algebra. We think this is the most elementary proof of this important result in the algebraic topology so far.